Christine went shopping and bought each of her nine nephews a gift, either a video costing $14.95 or a CD costing $16.88. She spent $138.41 on gifts. How many videos and CDs did she buy?
Answer by checkley77(12569)

v+c=9 or v=9-c
14.95v+16.88c=138.41
14.95(9-c)+16.88c=138.41
134.55-14.95c+16.88c=138.41
1.93c=138.41-134.55
1.93c=3.86
c=3.86/1.93
c=2 cds were bought.
v=9-2=7 video were bought.
Proof:
14.95*7+16.88*2=138.41
104.65+33.76=138.41
138.41=134.41